The Hydrogen Atom: The electron with charge -1.6 X 10 -19 C and proton with charge 1.6 X 10 -19 C of a hydrogen atom are separated by a distance of approximately 5.3 x 10 -11 m. Find the magnitude of the electric force between the two particles. 


1
Expert's answer
2022-03-04T08:31:44-0500
"F=k\frac{q_1q_2}{r^2}"

"F=9*10^9*\frac{1.6*10^{-19}*1.6*10^{-19}}{(5.3*10^{-11})^2}=8.2*10^{-8}\:\rm N"